OpenStudy (anonymous):

How did the lower and upper limits change? Here is the original equation: Integral of ( sin^7 theta ( 1-sin^2 theta ) cos theta ) dtheta FROM 0 to Pi/2 Let u = sin theta, du = cos theta dtheta The equation then becomes Integral of ( u^7 (1-u^2)^2 ) du From 0 to 1 Now how did it change from 0 to Pi/2, to 0 to 1?

OpenStudy (anonymous):

is very easy

OpenStudy (anonymous):

Tell me please :D

OpenStudy (turingtest):

\[\theta=0\implies u=\sin\theta=\sin0=0\]\[\theta=\frac\pi2\implies u=\sin\theta=\sin\frac\pi2=1\]

OpenStudy (anonymous):

yes

OpenStudy (anonymous):

Wow, thanks.